### v1

@@ Abstract
Context. Galaxy clusters are important cosmological probes that have helped to establish the $\Lambda\rm CDM$ paradigm as the standard model of cosmology. However, recent tensions between different types of high-accuracy data highlight the need for novel probes of the cosmological parameters. Such a probe is the turnaround density: the mass density on the scale where galaxies around a cluster join the Hubble flow.
Aims. To measure the turnaround density, one must locate the distance from the cluster center where turnaround occurs. Earlier work has shown that a turnaround radius can be readily identified in simulations by analyzing the 3D dark matter velocity field. However, measurements using realistic data face challenges due to projection effects. This study aims to assess the feasibility of measuring the turnaround radius using machine learning techniques applied to simulated observations of galaxy clusters.
Methods. We employ N-body simulations across various cosmologies to generate galaxy cluster projections. Utilizing Convolutional Neural Networks (CNNs), we assess the predictability of the turnaround radius based on galaxy line-of-sight velocity, number density, and mass profiles.
@@ 1 Introduction
Observations of the large scale structure in the Universe validate the $\Lambda\rm CDM$ model, with the exception of some apparent contradictions, referred to as “tensions”, particularly in relation to observations of the Hubble parameter and $\rm\sigma_{8}$ (see e.g., Bernal et al. 2016; Zhao et al. 2017; Joudaki et al. 2017; Hildebrandt et al. 2017; Riess et al. 2018, 2019; Motloch & Hu 2018; Planck Collaboration et al. 2020; Raveri & Hu 2019; Adhikari & Huterer 2019; Handley 2021; Di Valentino et al. 2020, 2021; Shah et al. 2021). Current observational evidence for the existence of a cosmological constant is based on the relation between the present-day values of the cosmological density parameters of matter, $\rm\Omega_{m}$ and dark energy, $\rm\Omega_{\Lambda}$, probed either by the cosmic microwave background or by observations of distant supernovae. Other cosmological datasets (cluster abundances, baryon acoustic oscillations) are also sensitive to $\rm\Omega_{m}$, but much less so to $\rm\Omega_{\Lambda}$. Recent work (Pavlidou et al. 2020; Pavlidou & Tomaras 2014; Korkidis et al. 2020, 2023; Tanoglidis et al. 2015, 2016) has demonstrated that the turnaround density ($\rm\rho_{ta}$) can be used to impose novel constraints on the cosmological parameters, highlighting that all structures share the same $\rm\rho_{ta}$ for the same redshift and cosmology.
@@
The turnaround scale (also known as the turnaround radius, $\rm R_{ta}$) marks the boundary where matter surrounding a collapsed cosmic structure transitions, on average, into the Hubble flow (Korkidis et al. 2020). Turnaround is not a singular event in a structure’s history. Instead, it is a continuous phenomenon observable around all structures at any given time. For example, around a big galaxy cluster, there is a region where material is actively accreting. However, beyond a certain distance, the expansion of the universe sweeps away other structures. The intersection of this accretion region with the Hubble flow occurs at the scale often referred to as the zero-velocity shell, the turnaround radius.
@@
To accurately determine $\rm\rho_{ta}$, it is essential to have knowledge of both $\rm R_{ta}$ and $\rm M_{ta}$ (the mass enclosed inside the turnaround radius). For very nearby objects it may be possible to reconstruct the turnaround mass and radius from the peculiar velocity field (e.g., Hoffman et al. 2018). However, in general, these methods do not scale with distance. The primary observational data that are widely available even for distant clusters are the distribution of galaxies as seen on the plane of the sky and their line of sight velocities. This limitation means that the turnaround radius cannot be reliably calculated by analyzing the velocity profiles directly; at a turnaround radius away from the center of the cluster as seen on the plane of the sky, the velocity component that transitions from positive to negative at turnaround is entirely on the plane of the sky, hence not observable. For the measurement to occur along the line of sight, distances of individual galaxies need to be measurable independently of their redshifts, and these are available only for nearby clusters. Consequently, researchers must rely on indirect methods or theoretical models to estimate the turnaround radius, such as the use of Zel’dovich sheets, as explored in Lee et al. (2015); Lee & Yepes (2016); Lee (2016); Lee & Li (2017); Lee (2018). Here, we explore a different approach, and test the feasibility of using machine learning tools directly on (simulated) galaxy position, mass and line-of-sight velocity data to identify the turnaround radius of a galaxy cluster.
@@
The use of machine learning methods, and in particular the subcategory of deep learning algorithms, i.e. Artificial Neural Networks (NNs), has experienced an exponential growth in astrophysical and cosmological studies over the recent years (Lahav 2023; Baron 2019; Carleo et al. 2019; Huertas-Company & Lanusse 2023), with several discussions on the effectiveness of these data driven methods (Lin et al. 2022). As per the universal approximation theorem, NNs—essentially compositions of weighted sums—are universal function approximators, capable of representing any continuous function in the limit of infinite parameters. Consequently, they serve as powerful tools for exploring data, with feature extraction capabilities embedded in their internal structure. For this reason, we adopt such an approach to investigate which types of observational datasets may encode information on the $\rm R_{ta}$ of galaxy clusters.
@@
In particular, we wish to establish whether there exists, in principle, a way of measuring the turnaround radius on the plane of the sky, using information on the distribution of galaxies around galaxy clusters, their masses, and their line-of-sight velocities. To this end, we employed N-body dark matter simulations to create projections of dark matter halo clusters (since baryonic effects are negligible at these scales, we use the terms “halo” and “galaxy” interchangeably) and measured their turnaround radii from their velocity profiles. We then applied machine learning techniques to establish a relationship between these projections and the accurate, numerically computed turnaround radii derived from the 3D information on the dark matter distribution and velocity field available in the simulations.
@@
The outline of this paper is as follows: In Sect. 2 we describe the set of N-body simulations used in this work and the characteristics of the resulting turnaround radii. In Sect. 3 we describe the methods used for the creation of the model that tries to predict the turnaround radius and in Sect. 4 we present the results of our analysis. We discuss these findings in Sect. 5.
@@ 2.1 N-body Simulations
In this work we use data from: (a) The MuldiDark Planck 2 (MDPL2) Simulation (Riebe et al. 2013). The MDPL2 simulation was performed with $(1{\rm Gpc}/h)^{3}$ box size, with about 57 billion particles and a cosmology consistent with Planck Collaboration et al. (2014). It was performed using the L-Gadget2 code. (b) The Virgo Intermediate Scale Simulations simulations ($\rm\Lambda CDM$, OCDM, SCDM) (Frenk et al. 2000; Thomas et al. 1998; Jenkins et al. 1998). These simulations contain $256^{3}$ particles each in a box of $(239.5\ {\rm Mpc}/h)^{3}$ and represent different cosmological models. The calculations were done using the external AP3M N-body code. For a more in depth description of the simulated boxes and the halo sample selection the reader is referred to Korkidis et al. 2023. For reference we include in Table 1 the main parameters of the 4 simulated boxes.
Similarly, we used the same cluster sample as in Korkidis et al. 2023 for the purposes of generating our mock cluster catalog, consisting of the largest $\sim 3200$ and $\sim 1000$ halos of the MDPL2 and Virgo simulations respectively. For each cluster in this sample, we calculated the turnaround radius kinematically by measuring the mean velocity of dark matter particles in spherical shells as described by Korkidis et al. (2020). The distribution of the turnaround radii values for the MDPL2 and Virgo $\Lambda$CDM simulations at $\rm z=0$, along with some 1-point statistics, can be seen in Fig. 1.
Table 1: Simulation box size in $h^{-1}$ Mpc, number of particles, particle mass in $h^{-1}\rm M_{\odot}$, Force resolution in units $h^{-1}$kpc and cosmological parameters. All simulations are from dark-matter–only runs.
MDPL2 1000 $3840^{3}$ $1.51\times 10^{9}$ $0.307115$ $0.048206$ $0.692885$ 0.6777
Virgo $\rm\Lambda CDM$ $239.5$ $256^{3}$ $6.86\times 10^{10}$ $0.3$ $0.7$ 0.70
Virgo SCDM $239.5$ $256^{3}$ $22.7\times 10^{10}$ $1.0$ $0.0$ $0.50$
Virgo OCDM $239.5$ $256^{3}$ $6.86\times 10^{10}$ $0.3$ $0.0$ $0.70$
@@ 2.2 Mock projections of Galaxy Clusters
Next, we projected all galaxy-sized halos ($\rm M_{200}\geq 10^{12}M_{\odot}$) around each cluster, onto each of the cluster’s three principal planes, retaining information on their positions and velocities perpendicular to the planes. These velocities would be equivalent to the line-of-sight velocities as observed by an observer situated far from the cluster. Employing three projections augmented our sample size by a factor of three, a crucial benefit for the effective training of our model. The projected area for each plane was chosen to be a square region of $25\ \rm Mpc$ on each side, centered on the respective cluster. This dimension was selected to ensure it would encompass the largest turnaround radius identified in our sample.
Lastly, the projection depth for the galaxies was informed by the varying uncertainties associated with different types of galaxy redshift surveys, specifically those employing either photometric or spectroscopic redshift measurements. Consequently, we created two distinct projected samples to reflect these different levels of redshift accuracy. In practical terms, this led to the imposition of a velocity cut on the velocities perpendicular to the projection plane, set at approximately $10^{3}\rm km/s$, and $10^{4}\rm km/s$, for spectroscopic and photometric redshift surveys, respectively.
Refer to captionRefer to caption
Figure 1: Histogram of the turnaround radii $\rm R_{ta}$ values from MDPL2 (upper panel) and Virgo $\rm\Lambda CDM$ (bottom panel). The difference between the two distributions reflects the difference between box sizes and resolutions.
All of the aforementioned steps were carried out at $z=0$. For the MDPL2 (Virgo $\rm\Lambda CDM$) data we additionally considered another 2 redshift snapshots, namely $z=0.49$ and $z=1.03$ ($z=0.5$ and $z=1.0$). In the histograms of Fig. 7 we can see the distribution of the $\rm R_{ta}$ values from each of the additional MDPL2 redshift snapshots (left panel) and from each of the non-$\rm\Lambda CDM$ Virgo simulations at $z=0$ (right panel). The different location of peaks in different redshifts and different cosmologies is a direct indication that $\rm\rho_{ta}$ is a function of redshift and cosmological parameters.
@@ 3.1 Post-processing
As mentioned in Sect. 2, our data consists of projected regions around cluster-sized dark matter halos, incorporating the details of each halo’s coordinates, virial mass and line-of-sight velocity. This format is not suitable for our analysis. Hence, in order to feed the data to the algorithm we converted them into images as 2D histograms. For the histograms we used appropriate weights to create 3 types of images: mass column density, number column density and mean line-of-sight velocity in each bin, in our case. We have not considered observational errors in each “observed” quantity. The reason for not adding noise is that we want to provide the model with the most favorable conditions possible. Should our present analysis fail to identify the turnaround scale, it is then certain that it will also fail in noisy datasets. However, if we succeed in identifying the turnaround scale, clearly further analysis will be needed to quantify the effect of noise, before proceeding to the analysis of real data.
For the first two cases (mass column density and number column density), the transformation does not change the nature of the data since we are merely normalizing by dividing by the area of each pixel (constant) to acquire the densities. For example for the mass column density we have:
@@
For the third case, by taking the mean velocity, some of the pixel values may change relative to one another in comparison with just binning the velocities, since we are dividing each pixel with a different number of halos, i.e
@@
where $\rm N_{i}$ the number of halos in pixel i. However, we discovered through several tests that this contributes positively to the outcome of our model. In addition to the original 3 types of images we also considered a $\rm 4^{th}$ type: the velocity dispersion in each bin, as a more promising (in terms of NN performance) and realistically observed quantity.
Different resolutions of the images were tested; we chose $25\times 25$ pixel images as optimal, both in terms of accuracy and computational efficiency. Higher resolution images have too many zero-valued pixels which appears to confuse the algorithm. Changes in the network’s structure with the use of l1 and l2 regularizations, which constrain the synaptic weights of the network, were not found to increase the high-resolution model’s performance. Similarly, changes in the nature of the projected images by scaling the data (min-max scaling, standardization) or by exchanging the 0-valued pixels with the mean or the median of the training set, produced unsatisfactory results. The issue was mitigated solely by the resolution choice.
We can see from Fig. 1 that the turnaround radius data are not uniform but rather they present a Gaussian distribution-like form (descending towards the right side due to the halo mass function and towards the left due to our sampling). This could potentially result in the algorithm only predicting values close to the peak, resulting in a problem usually referred to as mode collapse (Lin et al. 2022). We tried to transform the $\rm R_{ta}$ data so that the distribution of the values is less likely to affect the model. Namely we tried to pass $\rm R_{ta}$ through an invertible function to transform the data into a uniform distribution. The results showed that the original data were more successful and our final models are not affected by this issue.
Commonly, for each case, 80% of the simulation data were used for the training and validation and 20% were used for testing the trained model.

### v2

@@ Abstract
Context. Galaxy clusters are important cosmological probes that have helped to establish the $\Lambda$ cold dark matter paradigm as the standard model of cosmology. However, recent tensions between different types of high-accuracy data highlight the need for novel probes of the cosmological parameters. Such a probe is the turnaround density: the mass density on the scale where galaxies around a cluster join the Hubble flow. To measure the turnaround density, one must locate the distance from the cluster center where turnaround occurs. Earlier work has shown that a turnaround radius can be readily identified in simulations by analyzing the 3D dark matter velocity field. However, measurements using realistic data face challenges due to projection effects.
Aims. This study aims to assess the feasibility of measuring the turnaround radius using machine learning techniques applied to simulated idealized observations of galaxy clusters.
Methods. We employed N-body simulations across various cosmologies to generate galaxy cluster projections. Utilizing convolutional neural networks, we assessed the predictability of the turnaround radius based on galaxy line-of-sight velocity, number density, and mass profiles.
@@
Conclusions. Single-cluster turnaround radius inference from projected observables seems to be highly challenging. Future progress is likely to require statistical approaches, especially stacking, to exploit cosmological information encoded at turnaround scales.
@@ 1 Introduction
Table 1: Simulation box size in $h^{-1}$ Mpc, number of particles, particle mass in $h^{-1}\rm M_{\odot}$, force resolution in units of $h^{-1}$kpc and cosmological parameters. [^1] All simulations are from dark-matter–only runs.
MDPL2 1000 $3840^{3}$ $1.51\times 10^{9}$ $0.307115$ $0.048206$ $0.692885$ 0.6777
Virgo $\rm\Lambda CDM$ $239.5$ $256^{3}$ $6.86\times 10^{10}$ $0.3$ - $0.7$ 0.70
Virgo SCDM $239.5$ $56^{3} 22$$2.7\times 10^{10}$ $1.0$ - $0.0$ $0.50$
Virgo OCDM $239.5$ $256^{3}$ $6.86\times 10^{10}$ $0.3$ - $0.0$ $0.70$
@@
Observations of the large-scale structure in the Universe validate the $\Lambda$ cold dark matter ($\Lambda\rm CDM$) model, with the exception of some apparent contradictions, referred to as “tensions,” particularly in relation to observations of the Hubble parameter and $\rm\sigma_{8}$ (see e.g., Bernal et al. 2016; Zhao et al. 2017; Joudaki et al. 2017; Hildebrandt et al. 2017; Riess et al. 2018, 2019; Motloch & Hu 2018; Planck Collaboration et al. 2020; Raveri & Hu 2019; Adhikari & Huterer 2019; Handley 2021; Di Valentino et al. 2020, 2021; Shah et al. 2021; Vagnozzi 2023). Current observational evidence for the existence of a cosmological constant is based on the relation between the present-day values of the cosmological density parameters of matter, $\rm\Omega_{m}$ and dark energy, $\rm\Omega_{\Lambda}$, probed either by the cosmic microwave background or by observations of distant supernovae. Other cosmological datasets (cluster abundances, baryon acoustic oscillations) are also sensitive to $\rm\Omega_{m}$, but much less so to $\rm\Omega_{\Lambda}$. Recent work (Pavlidou et al. 2020; Pavlidou & Tomaras 2014; Korkidis et al. 2020, 2023; Tanoglidis et al. 2015, 2016) has demonstrated that the turnaround density ($\rm\rho_{ta}$) can be used to impose novel constraints on the cosmological parameters, highlighting that all structures share the same $\rm\rho_{ta}$ for the same redshift and cosmology.
@@
The turnaround scale, also known as the turnaround radius ($\rm R_{ta}$), marks the boundary where matter surrounding a collapsed cosmic structure transitions, on average, into the Hubble flow (Korkidis et al. 2020). Turnaround is not a singular event in a structure’s history. Instead, it is a continuous phenomenon observable around all structures at any given time. For example, around a big galaxy cluster, there is a region where material is actively accreting. However, beyond a certain distance, the expansion of the Universe sweeps away other structures. The intersection of this accretion region with the Hubble flow occurs at the scale often referred to as the zero-velocity shell, the turnaround radius.
@@
To accurately determine $\rm\rho_{ta}$, it is essential to know both $\rm R_{ta}$ and $\rm M_{ta}$ (the mass enclosed inside the turnaround radius). For very nearby objects, it may be possible to reconstruct the turnaround mass and radius from the peculiar velocity field (e.g., Hoffman et al. 2018). However, in general, these methods do not scale with distance. The primary observational data that are widely available, even for distant clusters, are the distribution of galaxies as seen on the plane of the sky and their line-of-sight velocities. This limitation means that the turnaround radius cannot be reliably calculated by analyzing the velocity profiles directly; at a turnaround radius away from the center of the cluster as seen on the plane of the sky, the velocity component that transitions from positive to negative at turnaround is entirely on the plane of the sky, hence not observable. For the measurement to occur along the line of sight, distances of individual galaxies need to be measurable independently of their redshifts, and these are available only for nearby clusters. Consequently, researchers must rely on indirect methods or theoretical models to estimate the turnaround radius, such as the use of Zel’dovich sheets, as explored in Lee et al. (2015); Lee & Yepes (2016); Lee (2016); Lee & Li (2017); Lee (2018). Here, we explore a different approach and test the feasibility of using machine learning tools directly on (simulated) galaxy position, mass, and line-of-sight velocity data to identify the turnaround radius of a galaxy cluster.
@@
The use of machine learning methods, and in particular the subcategory of deep learning algorithms, that is, artificial neural networks (NNs), has experienced an exponential growth in astrophysical and cosmological studies over the recent years (Lahav 2023; Baron 2019; Carleo et al. 2019; Huertas-Company & Lanusse 2023), with several discussions on the effectiveness of these data driven methods (Lin et al. 2022). As per the universal approximation theorem, NNs—essentially compositions of weighted sums—are universal function approximators, capable of representing any continuous function in the limit of infinite parameters. Consequently, they serve as powerful tools for exploring data, with feature extraction capabilities embedded in their internal structure. For this reason, we adopt such an approach to investigate which types of observational datasets may encode information on the $\rm R_{ta}$ of galaxy clusters.
@@
In particular, we wish to establish whether there exists in principle a way of measuring the turnaround radius on the plane of the sky, using information on the distribution of galaxies around galaxy clusters, their masses, and their line-of-sight velocities. To this end, we employed N-body dark matter simulations to create projections of dense dark matter halo regions (since baryonic effects are negligible at these scales, we use the terms “halo” and “galaxy” interchangeably) and measured their turnaround radii from their velocity profiles. For all purposes of this work, we use the term “cluster” to refer to these overdense regions. We then applied machine learning techniques to establish a relationship between these projections and the accurate, numerically computed turnaround radii derived from the 3D information on the dark matter distribution and velocity field available in the simulations.
@@
The outline of this paper is as follows: In Sect. 2 we describe the set of N-body simulations used in this work and the characteristics of the resulting turnaround radii. In Sect. 3 we describe the methods used for the creation of the model that aims to predict the turnaround radius, and in Sect. 4 we present the results of our analysis. We discuss these findings and conclude in Sect. 5.
@@ 2.1 N-body simulations
In this work, we used data from: (a) The MultiDark Planck 2 (MDPL2) Simulation (Riebe et al. 2013). The MDPL2 simulation was performed with $(1{\rm Gpc}/h)^{3}$ box size, with about 57 billion particles and a cosmology consistent with Planck Collaboration et al. (2014). It was performed using the L-Gadget2 code. (b) The Virgo Intermediate Scale Simulations of the $\rm\Lambda CDM$, open cold dark matter (OCDM), and standard cold dark matter (SCDM) models (Frenk et al. 2000; Thomas et al. 1998; Jenkins et al. 1998). These simulations contain $256^{3}$ particles each in a box of $(239.5\ {\rm Mpc}/h)^{3}$ and represent different cosmological models. The calculations were done using the external AP3M N-body code. For a more in-depth description of the simulated boxes and the halo sample selection, the reader is referred to Korkidis et al. (2023). For reference, we include in Table 1 the main parameters of the four simulated boxes.
Similarly, we used the same cluster sample as in Korkidis et al. (2023) for the purposes of generating our mock cluster catalog, corresponding to the largest $\sim 3200$ and $\sim 1000$ halos of the MDPL2 and Virgo simulations, respectively222with a small percentage of the MDPL2 resulting from the uniform secondary sampling of Korkidis et al. (2023) of logarithmically spaced bins at lower masses., (not making a distinction between halos and sub-halos as it is irrelevant in observations). For each cluster in this sample, we calculated the turnaround radius kinematically by measuring the mean velocity of dark matter particles in spherical shells as described by Korkidis et al. (2020), around the central (largest) halo.
Figure 1: Histogram of the turnaround radii $\rm R_{ta}$ values from MDPL2 (black) and Virgo $\rm\Lambda CDM$ (cyan). The difference between the two distributions reflects the difference between box sizes and resolutions. The mean ($\rm\mu$) and standard deviation ($\rm\sigma$) are shown on the top right, while, for the merged data $\rm(\mu,\sigma)=(7.8,2.5)$. Additional distributions can be found in Fig. 7 of the Appendix.
[…]
@@ 2.2 Mock projections of galaxy clusters
Next, we projected all galaxy-sized halos, $\rm M_{200}\geq 10^{12}M_{\odot}$ (corresponding to a stellar mass of $\rm M_{\star}\gtrsim 10^{10.5}M_{\odot}$ for $z=0$ depending on the model, see e.g., Girelli et al. 2020), around each cluster, onto each of the cluster’s three principal planes, retaining information on their positions, and velocities perpendicular to the planes. These velocities would be equivalent to the line-of-sight velocities as observed by an observer situated far from the cluster. Employing three projections augmented our sample size by a factor of three, a crucial benefit for the effective training of our model. The projected area for each plane was chosen to be a square region of $50\ \rm Mpc$ on each side, centered on the respective largest halo. This dimension was selected to ensure that it would encompass the largest turnaround radius identified in our sample.
Lastly, the projection depth for the galaxies was informed by the varying uncertainties associated with different types of galaxy redshift surveys, specifically those employing either photometric or spectroscopic redshift measurements. Consequently, we created two distinct projected samples to reflect these different levels of redshift accuracy. In practical terms, this led to the imposition of a velocity cut on the velocities perpendicular to the projection plane, set at approximately $10^{3}\rm km/s$ and $10^{4}\rm km/s$, for spectroscopic and photometric redshift surveys, respectively. While the latter does not fully capture photometric redshift uncertainties, it is sufficient for the scope of this work; a more detailed treatment is left for future studies.
All of the aforementioned steps were carried out at $z=0$. For the MDPL2 (Virgo $\rm\Lambda CDM$) data, we additionally considered another two redshift snapshots, namely $z=0.49$ and $z=1.03$ ($z=0.5$ and $z=1.0$). Throughout the paper, $z=0$ is implied when redshift is not explicitly stated. In the histograms of Fig. 7, we can see the distribution of the $\rm R_{ta}$ values from each of the additional MDPL2 redshift snapshots (left panel) and from each of the non-$\rm\Lambda CDM$ Virgo simulations at $z=0$ (right panel). The different location of peaks in different redshifts and different cosmologies is a direct indication that $\rm\rho_{ta}$ is a function of redshift and cosmological parameters.
@@ 3.1 Post-processing
As mentioned in Sect. 2, our data consists of projected regions around large dark matter halos, incorporating the details of each surrounding halo’s coordinates, virial mass, and line-of-sight velocity. This format was not suitable for our analysis (see also Sect. 3.2). Hence, in order to feed the data to the algorithm, we converted them into images as 2D histograms. For the histograms, we used appropriate weights to create three types of images: mass column density, number column density, and mean line-of-sight velocity in each bin, in our case (also abbreviated as “mass,” “vel,” and “num” hereafter). We did not consider observational errors in each “observed” quantity. The reason for not adding noise is that we wanted to provide the model with the most favorable conditions possible. Should our analysis fail to identify the turnaround scale, it is then certain that it will also fail in noisy datasets. However, if we succeed in identifying the turnaround scale, clearly, further analysis is needed to quantify the effect of noise before proceeding to the analysis of real data.
For the first two cases (mass column density and number column density), the transformation did not change the nature of the data since we merely normalized by dividing by the area of each pixel (constant) to acquire the densities. For example, for the mass column density, we have that
@@
For the third case, by taking the mean velocity, some of the pixel values could change relative to one another in comparison with just binning the velocities, since we divided each pixel by a different number of halos, that is,
@@
where $\rm N_{i}$ is the number of halos in pixel i. However, we discovered through several tests that this contributes positively to the outcome of our model. In addition to the original three types of images, we also considered a fourth type: the velocity dispersion in each bin, as a more promising (in terms of NN performance) and realistically observed quantity.
Different resolutions of the images were tested; we chose $25\times 25$ pixel images as optimal, both in terms of accuracy and computational efficiency, which allowed for more effective hyperparameter tuning [^3] The effects of image resolution and smoothing on the CNN’s performance can be found in Appendix C.. Higher resolution images had too many 0-valued pixels, which appeared to confuse the algorithm and made training with higher resolution images harder to converge, since the data are naturally sparse. Changes in the network’s structure with the use of L1 and L2 regularizations, which constrain the synaptic weights of the network, were not found to increase the high-resolution model’s performance. Similarly, changes in the nature of the projected images by scaling the data (min-max scaling, standardization) or by exchanging the 0-valued pixels with the mean or the median of the training set, produced unsatisfactory results.
We also treated $\rm R_{ta}$ as a deterministic target variable, since a more detailed treatment of its uncertainty would obscure the results of our feasibility study. We saw from Fig. 1 that the turnaround radius data are not uniform but rather present a Gaussian distribution-like form (descending toward the right side due to the halo mass function and toward the left due to our sampling as explained in Sect. 2). This could potentially result in the algorithm only predicting values close to the peak, resulting in a problem usually referred to as mode collapse (Lin et al. 2022). We tried to transform the $\rm R_{ta}$ data so that the distribution of the values was less likely to affect the model. Namely, we tried to pass $\rm R_{ta}$ through an invertible function to transform the data into a uniform distribution. The results showed that the original data were more successful, and our final models were not affected by this issue.
To train a NN, the splitting of the data is necessary. Commonly, for each case, 80% of the simulation data were used for the training and validation (with a ratio of 4:1). The remaining 20% were used for testing the trained model.